# On the Duality Between Sharpness-Aware Minimization and Adversarial Training

### v1

@@ 5 Experiment
Table 5: Comparison of variants of SAM on CIFAR-10 dataset.
Method Config Natural AA. StAdv
Standard SGD 94.5 31.7 35.2
Adam 93.9 13.9 20.4
SAM $\rho=0.1$ 95.4 43.6 39.3
$\rho=0.2$ 95.5 48.1 44.2
$\rho=0.3$ 95.4 57.8 46.1
$\rho=0.4$ 94.7 51.8 54.9
$\rho=0.5$ 94.5 58.4 55.8
ESAM $\rho=0.1$ 95.4 41.0 37.7
$\rho=0.2$ 95.6 47.5 43.0
$\rho=0.3$ 95.2 53.1 48.2
$\rho=0.4$ 94.7 51.6 49.6
$\rho=0.5$ 94.4 53.6 49.3
ASAM $\rho=0.1$ 95.2 36.4 35.6
$\rho=0.2$ 95.3 39.0 34.3
$\rho=0.3$ 95.4 39.9 34.7
$\rho=0.4$ 95.5 40.4 35.6
$\rho=0.5$ 95.7 44.2 36.6
@@
Table 7: Comparison on Tiny ImageNet.
Config Natural AA. StAdv
SGD 57.4 2.5 3.4
SAM ($\rho=0.4$) 57.9 10.3 13.3
SAM ($\rho=0.5$) 57.7 10.4 12.7
$\ell_{\infty}$-AT($\epsilon=8/255$) 32.3 27.7 25.1
$\ell_{2}$-AT($\epsilon=128/255$) 41.6 32.8 30.2
5.1 Image Classification
5.1.1 Experimental settings
Datasets and backbones. We examine the robustness of SAM on CIFAR-{10,100} (Krizhevsky et al., 2009) and TinyImageNet (Chrabaszcz et al., 2017) datasets. We mainly conduct our experiment with PreActResNet-18 (PRN-18) (He et al., 2016b). To demonstrate the scalability of SAM, we also include the Wider ResNet (WRN-28-10) (Zagoruyko and Komodakis, 2016) and the vision transformer architecture (Data-Efficient Image Transformer, DeiT)(Touvron et al., 2021).
Robustness evaluations. We assess the model’s resistance through a multifaceted approach to thoroughly explore its robustness against different threat models. We apply the Fast Gradient Sign Method (FGSM(Goodfellow et al., 2015) with a perturbation bound of $\ell_{\infty}$-$\epsilon=1/255$, and a 10-step Projected Gradient Descent (PGD(Madry et al., 2017) attack under $\ell_{\infty}$ and $\ell_{2}$ norms with bounds of $\epsilon=1/255$ and $32/255$, respectively. Moreover, we incorporate the AutoAttack (AA.) (Croce and Hein, 2021) to ensure a reliable evaluation. Besides, we also consider other popular attacks like StAdv (Xiao et al., 2018), FAB (Croce and Hein, 2020a) (10 steps under $\ell_{2}$-$\epsilon=32/255$). For black-box assessment, we utilize Pixle (Pomponi et al., 2022), a potent, pixel-rearranging black-box attack, constrained to a maximum of five iterations. The attacks are mostly implemented by the torchattacks (Kim, 2020) framework to ensure a reliable assessment. We ran all experiments three times independently and reported the average result. We omit the standard deviations since they are small (less than 0.5%) and do not affect our claims.
5.1.2 Experimental results
Comparison with standard training. As shown in Tables 2, 3 and 7, all models trained with SAM exhibit significantly better natural accuracy and robustness compared to those trained with standard training (ST). In particular, higher robustness is achieved by using larger values of $\rho$ with SAM. Taking the CIFAR-100 dataset as an example, the model trained with $\rho=0.4$ demonstrates even multiple robust accuracy than ST, and its natural accuracy is still higher than that of ST. Compared to the improvement in natural accuracy (approximately 2%), the increase of robustness is more significant (more than 10% in average). Moreover, on large datasets like Tiny-ImageNet, SAM still surpasses ST by a large margin. Therefore, we conclude that SAM with a relatively larger weight perturbation bound $\rho$ is a promising technique for enhancing adversarial robustness without sacrificing natural accuracy. Notably, our results are consistent with concurrent work showing that the robustness of SGD is better than Adam (Ma et al., 2023).
@@ 5.1.2 Experimental results
Variants of SAM. As shown in Table 5, SAM and ESAM generally outperform standard optimizers (SGD and Adam) in terms of robustness, demonstrating the effectiveness of adaptive momentum for adversarial defense. Notably, ESAM achieved the highest robustness against StAdv attacks (49.6% at $\rho=0.4$), while SAM yielded the best robustness against AA attacks (58.4% at $\rho=0.5$). These results highlight the robustness of various SAMs against threats.
@@ 5.2 Semantic Segmentation
Table 9: Comparing SAM and AT for semantic segmentation on VOC2012 dataset, mIoU on the val set is used for evaluation.
@@
Experimental settings. We train DeepLabv3 (Chen et al., 2017) with ResNet-50 (He et al., 2016a) backbone on the Stanford background dataset (Gould et al., ) with Cross-Entropy loss for 30000 iterations. The backbone is randomly initialized. And we use DeepLabv3 with mobilenetv2 (Sandler et al., 2018) backbone for experiments on VOC2012 dataset (Everingham et al., ), and the backbone is initialized by the weights pretrained on ImageNet. The learning rate is initialized as 0.01 with a polynomial decay scheme. For the optimizer, the weight decay is set to 1e-4, and the momentum is set to 0.9. We use the mean Intersection over Union (mIoU) (Everingham et al., 2015) to evaluate the segmentation results, where the IoU is calculated for each class at the pixel level as
${\rm IoU}=\frac{{\rm TP}}{{\rm TP}+{\rm FN}+{\rm FP}}\times 100\%,$ (15)
where TP represents true positive, FN is false negative and FP is false positive. The mIoU is then the mean value across all the classes in the dataset. For SAM, we select the perturbation hyper-parameter $\rho$ from the range $\{0.01,0.02,0.03,0.04,0.05\}$.
Experimental results. The experiment results are summarized in Table 8 and Table 9, where we use a 10-step PGD attack under $\ell_{\infty}$-$\epsilon=1/255$ and $\ell_{2}$-$\epsilon=1$ for robustness evaluation, respectively. As shown in the table, models trained using SAM all achieved better segmentation performance than ST and significantly better than AT. For robustness evaluation, all models trained by SAM exhibit better robustness than by ST, showcasing the generalization ability of SAM in various visual tasks. During robustness evaluation, the decrease in mIoU of segmentation tasks is not as significant as the decrease in accuracy of image classification tasks. This may be because the segmentation task is a highly interpretive task, in which the input and output results of the model are in the same dimension. In addition, many low-level features such as contour, texture, brightness, and good continuation also play an important role in the segmentation results (Ren and Malik, 2003; Zhu et al., 2021), which may make segmentation models inherently more robust than classification models.
@@ 5.3 Text Classification
Language models are also shown to be vulnerable against adversarial attacks (Morris et al., 2020; Wei et al., 2024). While adversarial attack and defense on Large Language Models (LLMs) have become emergent research topics recently (Zou et al., 2023; Dong et al., 2023; Wei et al., 2023c; Mo et al., 2024; Piet et al., 2023), it is worth noting that adversarial training on LLMs may not be helpful and practical (Jain et al., 2023), further underscoring the potential of deploying SAM a defense for language models.
Table 10: Natural accuracy and attack success rate (ASR) comparison between SAM and Standard Training (ST) for text sequence classification on Rotten Tomatoes Dataset.
@@
Accuracy ASR ($\downarrow$) ASR ($\downarrow$)
ST 84.4 22.7 84.1
@@
Experimental settings. In this part, we explore the effectiveness of SAM in enhancing the robustness of text sequence classification models against adversarial attacks. Utilizing the Rotten Tomatoes dataset (Pang and Lee, 2005) for sentiment analysis, our experiments provide insight into the resilience of these models when confronted with sophisticated word-level adversarial attacks. Specifically, we consider two popular attacks: the Improved Genetic Algorithm (IGA) (Wang et al., 2021) and the Particle Swarm Optimization (PSO) (Zang et al., 2020), both of which are officially supported by the TextAttack framework (Morris et al., 2020). For the model under test, we selected the widely recognized distilbert-base-uncased model from Hugging Face’s transformers library (Sanh et al., 2019), fine-tuning it for three epochs on the target dataset. The optimization for ST was performed using the AdamW optimizer with a learning rate of $5\times 10^{-5}$, and the SAM optimization strategy was applied with the same base learning rate for AdamW but incorporated a $\rho$ value of 0.05.
[…]
These findings suggest that incorporating sharpness-aware optimization techniques, such as SAM, can substantially improve the robustness of text classification models against increasingly sophisticated adversarial attacks without compromising natural language processing performance. This study contributes to the growing body of knowledge on defensive strategies in NLP, offering a promising avenue for future research and application to secure AI models against adversarial threats.
@@ 6 Conclusion
In this paper, we show that using Sharpness-Aware Minimization (SAM) alone can improve adversarial robustness, and reveal the duality relationship between SAM and Adversarial Training (AT). We empirically and theoretically demonstrate that both SAM and AT add perturbations to features to achieve better robust generalization ability. In addition, we illustrate that SAM can improve robustness without sacrificing natural performance, while AT inevitably hurts natural generalization. We further conduct experiments for image classification, semantic segmentation, and text classification on different datasets and various architectures, demonstrating the scalability and practicalness of using SAM to enhance robustness without sacrificing accuracy on clean data. Finally, we suggest that SAM can serve as a lightweight substitute for AT under certain requirements.
Impact Statements
@@ Acknowledgement
This work was sponsored by the Beijing Natural Science Foundation’s Undergraduate Initiating Research Program (Grant No. QY23041) and the National Natural Science Foundation of China (Grant No. 62172019).
@@ Appendix A Proofs
Lemma (Tsipras et al., 2018) In the presented model and data distribution, given $w_{1}$, the optimal solution for the optimization objective will assign equal weight to all the non-robust features, i.e. $w_{2}=w_{2}=\cdots=w_{n+1}$.
@@
Combining this with (32), we reach a contradiction.
@@
Since we only focus on $h$, we consider omitting the $o(h)$ terms in the calculation. To be more specific, $o(w^{2})$ term in the integral symbol $\int_{-h}^{0}$ can be omitted, and $o(w)$ or $o(h)$ term for $w$ and $h$ in the integral symbol $\int_{h}^{2\epsilon-2h}$ can also be omitted.The validity of abandoning these high order terms can be seen from the result of the calculation, which shows that $\frac{h-\epsilon}{h}\to 0$.
Combined with the proof in (33) and the definition of $w_{1}^{*}$, and abandoning the high order terms, we can calculate the right-hand side as follows.

### v2

@@ 5.1.2 Experimental results
Variants of SAM. As shown in Table 7, ASAM and ESAM generally outperform standard optimizers (SGD and Adam) in terms of robustness, demonstrating the effectiveness of adaptive momentum for adversarial defense. Notably, ESAM achieved the highest robustness against StAdv attacks (49.6% at $\rho=0.4$), while SAM yielded the best robustness against the AA attack (58.4% at $\rho=0.5$). These results highlight the robustness of various SAMs against adversarial threats.
Table 7: Comparison of variants of SAM on CIFAR-10 dataset.
Method Config Natural AA. StAdv
Standard SGD 94.5 31.7 35.2
Adam 93.9 13.9 20.4
SAM $\rho=0.3$ 95.4 57.8 46.1
$\rho=0.4$ 94.7 51.8 54.9
$\rho=0.5$ 94.5 58.4 55.8
ESAM $\rho=0.3$ 95.2 53.1 48.2
$\rho=0.4$ 94.7 51.6 49.6
$\rho=0.5$ 94.4 53.6 49.3
ASAM $\rho=0.3$ 95.4 39.9 34.7
$\rho=0.4$ 95.5 40.4 35.6
$\rho=0.5$ 95.7 44.2 36.6
Corruption Robustness. Unlike robustness against adversarial attacks, common corruption robustness captures a wider spectrum of realistic disturbances beyond artificially crafted perturbations, showcasing how models can handle varied and unexpected environmental changes. We evaluated SAM and AT on corrupted images in CIFAR-10C and calculated the average accuracy across 9 different common corruptions. We present these experimental results in Table 8, with detailed accuracy for each corruption type available in Table 13 in Appendix B.
Unlike in adversarial attack scenarios where SAM cannot outperform AT in terms of robustness, models trained with SAM have demonstrated significantly superior robustness compared to AT. As shown in the table, $\ell_{p}$-norm-AT performs even worse than standard training (15.67% for $\ell_{\infty}$-AT with $\epsilon=8/255$, compared to 34.91% for SGD). In contrast, SAM achieves similar results on corrupted images without significant performance degradation, showcasing the stability of SAM in realistic perturbation scenarios. Surprisingly, we found that variants of SAM can perform even better on corrupted images compared to standard training (with an accuracy increase of around 2%), highlighting the potential of the SAM method for addressing real-world robustness issues.
Table 8: Accuracy on common corruptions from CIFAR-10C.
Method Natural Corruption (Avg.)
SGD 94.5 34.91
Adam 93.9 29.60
SAM ($\rho=0.3$) 95.4 32.60
ASAM ($\rho=0.3$) 95.4 36.69
ESAM ($\rho=0.3$) 95.2 36.29
$\ell_{\infty}$-AT($\epsilon=8/255$) 84.5 15.67
$\ell_{2}$-AT($\epsilon=128/255$) 89.2 23.13
Comparison with AWP. Adversarial Weight Perturbation (AWP) (Wu et al., 2020) is an adversarial training method designed for better adversarial robustness by introducing a double-perturbation mechanism that adversarially perturbs both inputs and weights during training. Specifically, AWP can be regarded as a combination of SAM and AT. As AWP also leverages the weight perturbation paradigm, we provide a comparison with AWP and SAM to showcase that SAM still outperforms AWP in terms of natural accuracy.
The experimental results are shown in Table 9. Similar to the comparison between SAM and AT, we can see that although AWP outperforms both SAM and AT in terms of adversarial robustness, the natural accuracy with SAM is still higher than AWP even though weight perturbation was incorporated in it, indicating the importance of using natural data instead of adversarial examples for training to maintain the natural accuracy.
Table 9: Comparison of SAM, AT, and AWP on CIFAR-10.
Method Natural AA.
SAM ($\rho=0.1$) 95.4 43.6
SAM ($\rho=0.4$) 94.7 51.8
$\ell_{\infty}$-AT ($\epsilon=8/255$) 84.5 79.5
$\ell_{\infty}$-AWP ($\epsilon=8/255$) 82.0 80.1
$\ell_{2}$-AT ($\epsilon=128/255$) 89.2 84.8
$\ell_{2}$-AWP ($\epsilon=128/255$) 89.7 86.3
@@ 5.2 Semantic Segmentation
Experimental settings. We train DeepLabv3 (Chen et al., 2017) with randomly initialized ResNet-50 (He et al., 2016a) backbone on the Stanford background dataset (Gould et al., ) with Cross-Entropy loss for 30000 iterations. For experiments on VOC2012 dataset (Everingham et al., ), we use mobilenetv2 (Sandler et al., 2018) backbone, and the weights are initialized with pre-trained on ImageNet. The learning rate is initialized as 0.01 with a polynomial decay scheme. For the optimizer, the weight decay is set to 1e-4, and the momentum is set to 0.9. We use the mean Intersection over Union (mIoU) (Everingham et al., 2015) to evaluate the segmentation results, where the IoU is calculated for each class at the pixel level as
${\rm IoU}=\frac{{\rm TP}}{{\rm TP}+{\rm FN}+{\rm FP}}\times 100\%,$ (16)
where TP, FN and FP represent true positive, false negative and false positive, respectively. The mIoU is then the mean value across all the classes in the dataset. For SAM, we select $\rho$ from $\{0.01,0.02,0.03,0.04,0.05\}$.
@@
Table 11: Comparing SAM and AT for semantic segmentation on VOC2012 dataset, mIoU on the validation set is used for evaluation.
@@
Experimental results. The experiment results are summarized in Table 10 and Table 11, where we use a 10-step PGD attack under $\ell_{\infty}$-$\epsilon=1/255$ and $\ell_{2}$-$\epsilon=1$ for robustness evaluation, respectively. As shown in the table, models trained using SAM consistently achieve better segmentation performance than ST and significantly better than AT, and exhibit notably better robustness than ST, showcasing the robust generalization ability of SAM in various visual tasks. We also observed that the mIoU decrease of segmentation tasks is not as significant as the accuracy decrease of classification under adversarial attacks, which we assume is because the segmentation task is a highly interpretive task and many low-level features also play an important role in the segmentation results (Ren and Malik, 2003; Zhu et al., 2021), making segmentation models inherently more robust than classification models.
@@ 5.3 Text Classification
Language models are also shown to be vulnerable against adversarial attacks (Morris et al., 2020; Wei et al., 2024). While adversarial attack (Zou et al., 2023; Dong et al., 2023; Wei et al., 2023c; Zhang and Wei, 2024) and defense (Xie et al., 2023; Piet et al., 2023; Mo et al., 2024; Wang et al., 2024) on Large Language Models (LLMs) have become emergent research topics recently, it is worth noting that adversarial training on LLMs may not be helpful and practical (Jain et al., 2023), further underscoring the potential of deploying SAM a defense for language models.
Experimental settings. In this part, we explore the effectiveness of SAM in enhancing the robustness of text sequence classification models against adversarial attacks. Utilizing the Rotten Tomatoes dataset (Pang and Lee, 2005) for sentiment analysis, we study the word-level adversarial robustness of models trained with different methods. Specifically, we consider two popular attacks: the Improved Genetic Algorithm (IGA) (Wang et al., 2021) and the Particle Swarm Optimization (PSO) (Zang et al., 2020), both of which are officially supported by the TextAttack framework (Morris et al., 2020). For the model under test, we selected the widely recognized distilbert-base-uncased model from Hugging Face’s transformers library (Sanh et al., 2019), fine-tuning it for 3 epochs on the target dataset. The optimization for ST was performed using the AdamW optimizer with a learning rate of 5e-5, and the SAM optimization strategy was applied with the same base learning rate for AdamW and $\rho=0.05$. We also compare SAM with AT which incorporates adversarial examples identified in various adversarial attacks into the training dataset. Adversarial attacks such as IGA, PSO, and Textfooler (Jin et al., 2020) have been chosen for data augmentation purposes. The official implementation in TextAttack provides specific details for training in this regard.
[…]
Table 12: Accuracy and ASR comparison between SAM and ST for text sequence classification.
@@
Accuracy ($\uparrow$) ASR ($\downarrow$) ASR ($\downarrow$)
ST (AdamW) 84.4 22.7 84.1
@@
AT (w/IGA) 84.0 22.1 71.7
AT (w/PSO) 83.8 16.9 86.8
AT (w/Textfooler) 83.6 21.1 72.9
These findings suggest that incorporating SAM can notably improve the adversarial robustness of text classification models without compromising natural performance, offering a promising avenue for future research and application to secure AI models against adversarial threats.
@@ 6 Conclusion
In this paper, we reveal the duality relationship between Sharpness-Aware Minimization (SAM) and Adversarial Training (AT) and show that using SAM alone can improve adversarial robustness. We first intuitively illustrate that both SAM and AT can learn robust features by demonstrating the duality between weight and sample perturbations, and provide theoretical justifications to support these insights. We further conduct extensive experiments under various settings to show that SAM can improve robustness without sacrificing natural performance, while AT inevitably hurts natural generalization. These results uncover the scalability and practicality of using SAM to improve robustness without compromising accuracy on clean data. Based on this, we propose that SAM can be used as a lightweight alternative to AT when accuracy is a priority and improved robustness is preferred.
Impact Statement
@@ Acknowledgement
This work was sponsored by the Beijing Natural Science Foundation’s Undergraduate Initiating Research Program (Grant No. QY23041) and the National Natural Science Foundation of China (Grant No. 62172019, 623B2001, 82371112, 62394311).
@@ Appendix A Proofs
Lemma (Tsipras et al., 2018) In the presented model and data distribution, given $w_{1}$, the optimal solution for the optimization objective will assign equal weight to all the non-robust features, i.e. $w_{2}=w_{3}=\cdots=w_{n+1}$.
@@
Proof.
As $x_{1}$ has been chosen to be in ±$\pm 1$, the perturbation over $x_{1}$ has no influence on it, and we can just ignore it. Therefore, to attack the classifier by a bias $\boldsymbol{\delta}$, to make the accuracy as small as possible, an intuitive idea is to set $\boldsymbol{\delta}$ which minimizes the expectation of $x_{i}(i=2,\cdots,n+1),$ which made the standard accuracy smaller. In fact, the expected accuracy is monotonically increasing about each $\delta_{i}(i=2,\cdots,n+1).$ Thus, choosing $\delta=(0,-\epsilon,\cdots,-\epsilon)$ can be the best adversarial attack vector for any $w>0$. In this situation, this equals $x_{i}^{\prime}(i=2,\cdots,n+1)\sim\mathcal{N}(\eta-\epsilon,n).$
Thus, $R_{A}$ can be rewritten to be
$R_{A}$ $=\mathbb{E}_{\mathbf{x},y\sim D}\mathbb{P}(w_{1}x_{1}+\sum_{i=2}^{n+1}x_{i}-n\epsilon>0)$ (17)
$=p\Phi((w_{1}+(\eta-\epsilon)n)/\sqrt{n})+(1-p)\Phi((-w_{1}+(\eta-\epsilon)n)/\sqrt{n}),$
where $\Phi(\cdot)$ is the cumulative distribution function of a standard normal distribution.
Since $W_{R}$ is obviously a monotonic increasing function of $w_{1}$, we only need to prove that $RA$ in the above equation is a monotonic increasing function of $w_{1}$. To show this, we take a partial derivative of $w_{1}$ in $RA$:
$\frac{\partial RA}{\partial w_{1}}=p\exp({-(w_{1}+(\eta-\epsilon)n)^{2}/2n})/\sqrt{2\pi n}-(1-p)\exp({-(w_{1}-(\eta-\epsilon)n)^{2}/2n})/\sqrt{2\pi n}.$ (18)
Simplifying this, we get
$\frac{\partial R_{A}}{\partial w_{1}}=g(w_{1})\cdot(1-\frac{1-p}{p}\exp(2w_{1}(\eta-\epsilon))),$ (19)
where $g(w_{1})=\frac{p}{\sqrt{2\pi n}}\exp({-(w_{1}+(\eta-\epsilon)n)^{2}/2n})$ is positive . When $0<W_{R}<W_{R}^{AT}$, or equivalantly $0<w_{1}<\frac{\ln p-\ln(1-p)}{2(\eta-\epsilon)}$, we can easily know that $(1-\frac{1-p}{p}\exp(2w_{1}(\eta-\epsilon)))>0$, which implies that $\frac{\partial R_{A}}{\partial w_{1}}>0$, and further means that RA is a monotonic increasing function of $W_{R}$. ∎
A.2 Proof for Theorem 4.2
@@
Combining this with (LABEL:eq21), we reach a contradiction.
@@
Since we only focus on $h$, we consider omitting the $o(h)$ terms in the calculation. To be more specific, $o(w^{2})$ term in the integral symbol $\int_{-h}^{0}$ can be omitted, and $o(w)$ or $o(h)$ term for $w$ and $h$ in the integral symbol $\int_{h}^{2\epsilon-2h}$ can also be omitted. [^1] The validity of abandoning these high order terms can be seen from the result of the calculation, which shows that $\frac{h-\epsilon}{h}\to 0$.
Combined with the proof in (LABEL:eq22) and the definition of $w_{1}^{*}$, and abandoning the high order terms, we can calculate the right-hand side as follows.
@@ A.6 Proof for Theorem 5
Appendix B Detailed Experimental Results
Overall experimental results for general robust accuracy are shown in Table 13. Method Natural Brightness Fog Frost Gaussian Impulse Jpeg Shot Snow Speckle Accuracy Blur Noise Noise Noise SGD 94.5 63.7 26.3 30.8 19.2 40.6 32.4 26.5 41.9 32.8 Adam 93.9 54.8 21.9 25.4 15.9 35.7 28.1 21.5 38.0 25.1 SAM ($\rho=0.1$) 95.4 63.1 26.2 29.2 20.6 37.0 31.7 23.0 42.7 28.3 SAM ($\rho=0.2$) 95.5 61.5 20.2 27.5 16.6 40.2 27.2 26.1 38.1 30.2 SAM ($\rho=0.3$) 95.4 61.0 24.8 28.6 19.1 40.3 28.2 23.7 40.2 27.5 SAM ($\rho=0.4$) 94.7 59.5 22.4 31.8 17.0 37.7 30.0 29.7 41.7 35.3 ASAM ($\rho=0.3$) 95.4 68.9 28.7 34.0 18.2 41.9 34.0 25.9 45.9 32.7 ESAM ($\rho=0.3$) 95.2 64.3 28.1 32.9 20.9 39.3 32.8 30.0 44.2 34.1 AT ($\ell_{\infty}$-$\epsilon=\frac{8}{255}$) 84.5 16.1 12.5 11.6 15.5 18.2 16.7 17.3 16.2 16.9 AT ($\ell_{2}$-$\epsilon=\frac{128}{255}$) 89.2 25.8 14.6 18.7 21.5 25.3 25.4 26.3 25.2 25.4
Table 13: General robust accuracy on CIFAR-10C dataset.